# Open Materials Generation with Stochastic Interpolants

### v1

@@ Abstract
The discovery of new materials is essential for enabling technological advancements. Computational approaches for predicting novel materials must effectively learn the manifold of stable crystal structures within an infinite design space. We introduce Open Materials Generation (OMG), a unifying framework for the generative design and discovery of inorganic crystalline materials. OMG employs stochastic interpolants (SI) to bridge an arbitrary base distribution to the target distribution of inorganic crystals via a broad class of tunable stochastic processes, encompassing both diffusion models and flow matching as special cases. In this work, we adapt the SI framework by integrating an equivariant graph representation of crystal structures and extending it to account for periodic boundary conditions in unit cell representations. Additionally, we couple the SI flow over spatial coordinates and lattice vectors with discrete flow matching for atomic species. We benchmark OMG’s performance on two tasks: Crystal Structure Prediction (CSP) for specified compositions, and de novo generation (DNG) aimed at discovering stable, novel, and unique structures. In our ground-up implementation of OMG, we refine and extend both CSP and DNG metrics compared to previous works. OMG establishes a new state-of-the-art in generative modeling for materials discovery, outperforming purely flow-based and diffusion-based implementations. These results underscore the importance of designing flexible deep learning frameworks to accelerate progress in materials science.
@@ I Introduction
A core objective of materials science is the discovery of new synthesizable structures and compounds with the potential to meet critical societal demands. The development of new materials such as room-temperature superconductors [1], high-performance alloys with exceptional mechanical properties [2, 3, 4], advanced catalysts [5, 6], and materials for energy storage and generation [7, 8] hold the potential to drive technological revolutions.
@@
Still, efficiently sampling the manifold of stable materials structures under diverse constraints—such as composition and target properties—remains a major challenge. Traditional approaches to materials discovery have relied on first-principles electronic structure methods such as DFT—or more sophisticated theory, depending on the property [22, 23, 24]—which, while powerful and fairly accurate, are very computationally expensive. These methods include ab initio random structure searching (AIRSS) [25] or genetic algorithms for structure and phase prediction [26], both of which have successfully predicted new crystal structures and some of which have even been experimentally realized [27]. However, the high computational cost of these approaches has limited the scope and speed of material exploration, highlighting the need for cutting-edge ML techniques to significantly accelerate the discovery of stable inorganic crystalline materials.
@@ I.1 Related works
While these approaches have demonstrated that ML can push the boundaries of computational materials discovery, it remains uncertain whether score-based diffusion or flow-matching represent the definitive methodological frameworks for this problem. Furthermore, the extent to which the optimal approach depends on the training data remains an open question. Thus far, each new method has typically outperformed its predecessors across datasets.
@@ I.2 Our contribution
The work we present in this paper is the first implementation and extension of the stochastic interpolants (SIs) framework [33] for the modeling and generation of inorganic crystalline materials. SIs are a unifying framework that encompasses both flow-based and diffusion-based methods as specific instances, while offering a more general and flexible framework for generative modeling. In this context, SIs define a stochastic process that interpolates between pairs of samples from a known base distribution and a target distribution of inorganic crystals. By learning the drift terms of either ordinary differential equations (ODEs) or stochastic differential equations (SDEs), new samples can be generated by numerically integrating these equations. The flexibility of the SI framework stems from the ability to tailor the choice of interpolants, and the incorporation of an additional random latent variable, further enhancing its expressivity. With their many degrees of freedom, SIs thus provide an ideal framework for optimizing generative models for materials design.
We implement the SI framework in the open-source Open Materials Generation (OMG) package, released alongside this paper. OMG allows to train and benchmark models for two materials generation tasks: Crystal structure prediction (CSP) which only learns to generate atomic positions and lattice vectors for a given composition, and de novo generation (DNG) which learns to generate both crystal structure and composition to predict novel materials. We discover that optimizing interpolation schemes for different degrees of freedom—particularly fractional coordinates in the crystal representation—substantially improves performance across diverse datasets. As a result, our approach achieves a new state of the art—outperforming both DiffCSP [30] and FlowMM [32]—in CSP and DNG across all evaluated datasets for both existing, revised, and new performance measures.
@@ II.1 Diffusion models
A widely used approach in generative modeling uses diffusion models [34], which define a stochastic process that progressively transforms structured data into noise. A model is then trained to approximate the reverse process, enabling the generation of new samples. These models are often formulated in terms of SDEs, where the forward process follows a predefined diffusion dynamic.
@@ II.2 Conditional flow matching
Conditional flow matching (CFM) [38] is a generative modeling technique that learns a flow which transports samples from a base distribution at time $t=0$ to a target distribution at time $t=1$. This process defines a probability path that describes how samples are distributed at any intermediate time $t\in[0,1]$. The velocity field associated with this flow governs how individual samples evolve over time. CFM learns the velocity indirectly by constructing conditional vector fields that are known a priori. Once trained, samples drawn from the base distribution can be evolved numerically to generate new samples from the target distribution. Originally, CFM was formulated using Gaussian conditional probability paths, but Tong et al. [39] later extended this framework to allow for arbitrary probability paths and couplings between base and target distributions. A further extension, particularly relevant to physics and chemistry, is Riemannian flow matching (RFM), which generalizes CFM to Riemannian manifolds [40]. This allows in particular to use the flow-matching framework for systems with periodic boundary conditions as they appear in unit cell representations of inorganic crystals [32].
@@ III.1 Stochastic Interpolants
SIs provide a unifying mathematical framework for generative modeling, generalizing both SBDMs and CFM [33]. The SI $x(t,x_{0},x_{1},z)$ bridges the base distribution $\rho_{0}$ with a target distribution $\rho_{1}$ by learning a time-dependent drift $b^{\theta}(t,x)$. In this work, we focus on stochastic interpolants between $\rho_{0}$ and $\rho_{1}$ of the form:
@@
Here, $t\in[0,1]$ represents time and $(x_{0},x_{1})$ are paired samples drawn from $\rho_{0}$ and $\rho_{1}$, respectively. The random variable $z$ is drawn from a standard Gaussian $\mathcal{N}(0,\bm{I})$ independently of $x_{0}$ and $x_{1}$. The functional forms of $\alpha$, $\beta$, and $\gamma$ are flexible, subject to few few constraints (see Appendix A.2). The inclusion of the latent variable $\gamma(t)z$ allows sampling of an ensemble of paths around the mean interpolant $I(t,x)=\alpha(t)x_{0}+\beta(t)x_{1}$, and is theorized to improve generative modeling by promoting smoother and more regular learned flows [33].
For generative modeling, it is most important that the time-dependent density $\rho(t)$ of the stochastic process $x_{t}$ in Eq. 1 can also be realized either via deterministic sampling through an ODE (derived from a transport equation) or stochastic sampling through an SDE (derived from a Fokker–Planck equation) only requiring $x_{0}\sim\rho_{0}$ (see Appendix A.1). For both ODE- and SDE-based sampling, the required drift term $b^{\theta}(t,x):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is learned by minimizing the loss function
$\begin{split}\mathcal{L}_{b}(\theta)=&\,\mathbb{E}_{t,z,x_{0},x_{1}}\\ &\big{[}|b^{\theta}(t,x_{t})|^{2}-2\,\partial_{t}x(t,x_{0},x_{1},z)\cdot b^{\theta}(t,x_{t})\big{]},\end{split}$ (2)
where the expectation is taken independently over $t\sim\mathcal{U}(0,1)$ with $\mathcal{U}(0,1)$ the uniform distribution between 00 and $1$, $z\sim\mathcal{N}(0,\bm{I})$, $x_{0}\sim\rho_{0}$, and $x_{1}\sim\rho_{1}$. For SDE-based sampling, an additional denoiser $z^{\theta}(t,x):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ must be learned by minimizing an additional loss
@@
The drift term, along with the denoiser in the case of SDE-based sampling, enables the generation of samples from the target distribution [33]. For ODE-based sampling, $\gamma(t)=0$ is a possible choice. For SDE-based sampling, however, $\gamma(t)>0$ is required for all $t\in(0,1)$ (see Appendix A.1).
By appropriately selecting interpolation functions $\alpha$, $\beta$, $\gamma$ and choosing between deterministic (ODE) and stochastic (SDE) sampling schemes, the SI framework not only recovers CFM and SBDM as special cases but also enables the design of a broad class of novel generative models (see Appendix A.2 for examples). The strength of OMG’s SI implementation for materials discovery lies in its ability to tune both the interpolation and sampling schemes, as illustrated in Fig. 1 for a pair of structures sampled from $\rho_{0}$ and $\rho_{1}$. By systematically optimizing over this large design space, we achieve superior performance for CSP and DNG tasks across datasets, as discussed in Section V.
Figure 1: Visualization of the tunable components of the SI framework for bridging samples $x_{0}$ (gray particles) and $x_{1}$ (purple particles). (a) The choice of the interpolant changes the path of the time-dependent interpolation trajectory. (b) During inference, the learned drift term $b^{\theta}(t,x)$ and denoiser $z^{\theta}(t,x)$ generate new samples via standard ODE or SDE integration, here for a linear interpolant with $\gamma=\sqrt{0.07t(1-t)}$. (c) The inclusion of a latent variable $\gamma(t)z$ changes the interpolation path. (d) The function $\gamma(t,a)=\sqrt{at(1-t)}$ depends on $a$ that also influences the interpolation path.
@@ III.2 Crystal representation and generation
A crystalline material is defined by its idealized repeat unit, or unit cell, which encodes its periodicity. In the OMG representation, a unit cell is described by separating the material’s chemical composition—given by its atomic species $\bm{A}\in\mathbb{Z}_{>0}^{N}$, where $N$ is the number of atoms in the unit cell—from its structural representation—its fractional coordinates $\bm{X}\in[0,1)^{3\times N}$ with periodic boundaries and lattice vectors $\bm{L}\in\mathbb{R}^{3\times 3}$. During training, all three components $\left\{\bm{A},\bm{X},\bm{L}\right\}$ are considered simultaneously. We apply the SI framework only to the continuous structural representations $\left\{\bm{X},\bm{L}\right\}$ with loss functions defined in Eq. 2 and 3, and use discrete flow matching (DFM) on the chemical species $\bm{A}$ [41]. The number of atoms $N$ in the structure $x_{0}$ sampled from the base distribution $\rho_{0}$ is determined by the number of atoms in the corresponding structure $x_{1}$ sampled from the target distribution $\rho_{1}$.
@@ III.2.1 Atomic coordinates
We specify the base distribution for the fractional coordinates $\bm{x}\in[0,1)$ for all $\bm{x}\in\bm{X}$ via a uniform distribution (except for the score-based diffusion interpolant that requires a wrapped normal distribution $\rho_{0}(\bm{x})$ following the approach of Jiao et al. [30]; see Section IV.1). For learning fractional coordinates, we implement a variety of periodic interpolants that connect the base to the target data distributions. To satisfy periodic boundary conditions on the paths defined by the interpolants, we extend the SI framework to the surface of a four-dimensional torus in this paper. Reminiscent of RFM [40], the linear interpolant on the torus traverses a path equivalent to the shortest-path geodesic which is always [^1] The only exception being when two points are precisely half the box length apart. However, this case is not relevant for the given base distribution. well-defined. Other interpolants, however, are more complex. In order to uniquely define them, we always define the interpolation with respect to the shortest-path geodesic. That is, for interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with a periodic boundary at 0 and $1$, we first unwrap $\bm{x}_{1}$ to the periodic image $\bm{x}_{1}^{\prime}$ which has the shortest possible distance from $\bm{x}_{0}$. Following this, the interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}^{\prime}$ is computed given a choice of interpolant, and the traversed path is wrapped back into the boundary from 00 to $1$. This approach is illustrated in Appendix A.4.
@@ III.2.2 Lattice vectors
To construct the base distribution for the cell vectors $\bm{L}$, we follow Miller et al. [32] and use an informed log-normal distribution $\rho_{0}(\bm{L})$ which is fitted to the distribution of cell vectors in each target dataset. This leverages the flexibility in the SI framework to choose arbitrary base distributions in contrast to SBDM which necessarily requires a Gaussian base distribution.
@@ III.2.3 Atomic species
In doing this, we are able to directly construct the marginal rate matrix $R^{\theta}_{t}(\bm{a}_{t},i)$ for the CTMC that dictates the evolution of $\bm{a}_{t}$ at time $t$ to the next time step during generation (see Appendix A.5). It is important to note that the learned probability path is a function of the entire atomic configuration $\left\{\bm{A},\bm{X},\bm{L}\right\}$ which is necessary for the prediction of chemical composition from structure.

### v2

@@ Abstract
The discovery of new materials is essential for enabling technological advancements. Computational approaches for predicting novel materials must effectively learn the manifold of stable crystal structures within an infinite design space. We introduce Open Materials Generation (OMatG), a unifying framework for the generative design and discovery of inorganic crystalline materials. OMatG employs stochastic interpolants (SI) to bridge an arbitrary base distribution to the target distribution of inorganic crystals via a broad class of tunable stochastic processes, encompassing both diffusion models and flow matching as special cases. In this work, we adapt the SI framework by integrating an equivariant graph representation of crystal structures and extending it to account for periodic boundary conditions in unit cell representations. Additionally, we couple the SI flow over spatial coordinates and lattice vectors with discrete flow matching for atomic species. We benchmark OMatG’s performance on two tasks: Crystal Structure Prediction (CSP) for specified compositions, and de novo generation (DNG) aimed at discovering stable, novel, and unique structures. In our ground-up implementation of OMatG, we refine and extend both CSP and DNG metrics compared to previous works. OMatG establishes a new state of the art in generative modeling for materials discovery, outperforming purely flow-based and diffusion-based implementations. These results underscore the importance of designing flexible deep learning frameworks to accelerate progress in materials science. The OMatG code is available at https://github.com/FERMat-ML/OMatG.
@@ I Introduction
A core objective of materials science is the discovery of new synthesizable structures and compounds with the potential to meet critical societal demands. The development of new materials such as room-temperature superconductors [1], high-performance alloys with exceptional mechanical properties [2, 3, 4], advanced catalysts [5, 6], and materials for energy storage and generation [7, 8] holds the potential to drive technological revolutions.
@@
Still, efficiently sampling the manifold of stable materials structures under diverse constraints—such as composition and target properties—remains a major challenge. Traditional approaches to materials discovery have relied on first-principles electronic structure methods such as density functional theory (DFT) [^1] See Appendix A for a list of acronyms used throughout this paper.—or more sophisticated theory, depending on the property [22, 23, 24]—which, while powerful and fairly accurate, are very computationally expensive. These methods include ab initio random structure searching (AIRSS) [25] or genetic algorithms for structure and phase prediction [26], both of which have successfully predicted new crystal structures and some of which have even been experimentally realized [27]. However, the high computational cost of these approaches has limited the scope and speed of material exploration, highlighting the need for cutting-edge ML techniques to significantly accelerate the discovery of stable inorganic crystalline materials.
@@ I.1 Related Works
While these approaches have demonstrated that ML can push the boundaries of computational materials discovery, it remains uncertain whether score-based diffusion or flow-matching represents the definitive methodological frameworks for this problem. Furthermore, the extent to which the optimal approach depends on the training data remains an open question. Thus far, each new method has typically outperformed its predecessors across datasets.
@@ I.2 Our Contribution
The work we present in this paper is the first implementation and extension of the stochastic interpolants (SIs) framework [33] for the modeling and generation of inorganic crystalline materials. SIs are a unifying framework that encompasses both flow-matching and diffusion-based methods as specific instances, while offering a more general and flexible framework for generative modeling. In this context, SIs define a stochastic process that interpolates between pairs of samples from a known base distribution and a target distribution of inorganic crystals. By learning the velocity term of an ordinary differential equation (ODE) or the drift term of a stochastic differential equations (SDE), new samples can be generated by numerically integrating these equations. The flexibility of the SI framework stems from the ability to tailor the choice of interpolants, and the incorporation of an additional random latent variable, further enhancing its expressivity. With their rich parameterization, SIs thus provide an ideal framework for optimizing generative models for materials design.
We implement the SI framework in the open-source Open Materials Generation (OMatG) package, released alongside this paper. OMatG allows to train and benchmark models for two materials generation tasks: Crystal structure prediction (CSP) which only learns to generate atomic positions and lattice vectors for a given composition, and de novo generation (DNG) which learns to generate both crystal structure and composition to predict novel materials. We discover that optimizing interpolation schemes for different degrees of freedom of the crystal unit cell substantially improves performance across diverse datasets. As a result, our approach achieves a new state of the art—outperforming both DiffCSP [30] and FlowMM [32] in CSP and DNG, as well as MatterGen [31] in DNG—across all evaluated datasets under existing, revised, and new performance measures.
@@ II.1 Diffusion Models
A widely used approach in generative modeling uses diffusion models [34], which define a stochastic process that progressively transforms structured data into noise via a predefined diffusion dynamic. A model is then trained to approximate the reverse process, enabling the generation of new samples, typically by integrating a corresponding SDE.
@@ II.2 Conditional Flow Matching
Conditional flow matching (CFM) [38, 39, 40] is a generative modeling technique that learns a flow which transports samples from a base distribution at time $t=0$ to a target distribution at time $t=1$. This process defines a probability path that describes how samples are distributed at any intermediate time $t\in[0,1]$. The velocity field associated with this flow governs how individual samples evolve over time. CFM learns the velocity indirectly by constructing conditional vector fields that are known a priori. Once trained, samples drawn from the base distribution can be evolved numerically to generate new samples from the target distribution. Originally, CFM was formulated using Gaussian conditional probability paths, but Tong et al. [41] later extended this framework to allow for arbitrary probability paths and couplings between base and target distributions. A further extension, particularly relevant to physics and chemistry, is Riemannian flow matching (RFM), which generalizes CFM to Riemannian manifolds [42]. This allows in particular to use the flow-matching framework for systems with periodic boundary conditions as they appear in unit cell representations of inorganic crystals [32].
@@ III.1 Stochastic Interpolants
SIs provide a unifying mathematical framework for generative modeling, generalizing both SBDMs and CFM [33]. The SI $x(t,x_{0},x_{1},z)$ bridges the base distribution $\rho_{0}$ with a target distribution $\rho_{1}$ by learning a time-dependent map. In this work, we focus on stochastic interpolants of the form:
@@
Here, $t\in[0,1]$ represents time and $(x_{0},x_{1})$ are paired samples drawn from $\rho_{0}$ and $\rho_{1}$, respectively. The random variable $z$ is drawn from a standard Gaussian $\mathcal{N}(0,\bm{I})$ independently of $x_{0}$ and $x_{1}$. The functional forms of $\alpha$, $\beta$, and $\gamma$ are flexible, subject to few constraints (see Appendix B.2). The inclusion of the latent variable $\gamma(t)z$ allows sampling of an ensemble of paths around the mean interpolant $I(t,x)=\alpha(t)x_{0}+\beta(t)x_{1}$, and is theorized to improve generative modeling by promoting smoother and more regular learned flows [33].
The time-dependent density $\rho_{t}$ of the stochastic process $x_{t}$ in Eq. (1) can be realized either via deterministic sampling through an ODE (derived from a transport equation) or stochastic sampling through an SDE (derived from a Fokker–Planck equation) only requiring $x_{0}\sim\rho_{0}$ (see Appendix B.1). This enables generative modeling by evolving samples from a known base distribution $\rho_{0}$ to the target distribution $\rho_{1}$. For both ODE- and SDE-based sampling, the required velocity term $b^{\theta}(t,x)$ is learned by minimizing the loss function
$\begin{split}\mathcal{L}_{b}(\theta)=\,\mathbb{E}_{t,z,x_{0},x_{1}}\big{[}&|b^{\theta}(t,x_{t})|^{2}\\ &-2\,\partial_{t}x(t,x_{0},x_{1},z)\cdot b^{\theta}(t,x_{t})\big{]},\end{split}$ (2)
where the expectation is taken independently over $t\sim\mathcal{U}(0,1)$, the uniform distribution between 00 and $1$, $z\sim\mathcal{N}(0,\bm{I})$, $x_{0}\sim\rho_{0}$, and $x_{1}\sim\rho_{1}$. For SDE-based sampling, an additional denoiser $z^{\theta}(t,x)$ must be learned by minimizing an additional loss
@@
The velocity term, along with the denoiser in the case of SDE-based sampling, enables the generation of samples from the target distribution [33]. Note that minimizing with respect to these loss functions amounts to minimizing with respect to a mean-squared error loss function (see Appendix C.2). For ODE-based sampling, $\gamma(t)=0$ in the interpolant $x(t,x_{0},x_{1},z)$ is a possible choice. However, for SDE-based sampling, $\gamma(t)>0$ is required for all $t\in(0,1)$ (see Appendix B.1).
By appropriately selecting interpolation functions $\alpha$, $\beta$, $\gamma$ and choosing between deterministic (ODE) and stochastic (SDE) sampling schemes (see Appendix B.2 for examples), the SI framework not only recovers CFM and SBDM as special cases (see Appendix B.9) but also enables the design of a broad class of novel generative models. The strength of OMatG’s SI implementation for materials discovery lies in its ability to tune both the interpolation and sampling schemes, as illustrated in Fig. 1 for a pair of structures sampled from $\rho_{0}$ and $\rho_{1}$. By systematically optimizing over this large design space, we achieve superior performance for CSP and DNG tasks across datasets, as discussed in Section V.
Figure 1: Visualization of the tunable components of the SI framework for bridging samples $x_{0}$ (gray particles) and $x_{1}$ (purple particles). Interpolation paths are shown only for one pair of highlighted particles. (a) The choice of the interpolant changes the path of the time-dependent interpolation trajectory. (b) During inference, the learned velocity term $b^{\theta}(t,x)$ and denoiser $z^{\theta}(t,x)$ generate new samples via ODE or SDE integration, here for a linear interpolant with $\gamma=\sqrt{0.07t(1-t)}$. (c) The inclusion of a latent variable $\gamma(t)z$ changes the interpolation path. For SDE-based sampling, $\gamma(t)>0$ is required. (d) The function $\gamma(t,a)=\sqrt{at(1-t)}$ depends on $a$ that also influences the interpolation path.
@@ III.2 Crystal Representation and Generation
A crystalline material is defined by its idealized repeat unit, or unit cell, which encodes its periodicity. In the OMatG representation, a unit cell is described by separating the material’s chemical composition—given by its atomic species $\bm{A}\in\mathbb{Z}_{>0}^{N}$, where $N$ is the number of atoms in the unit cell—from its structural representation—its fractional coordinates $\bm{X}\in[0,1)^{3\times N}$ with periodic boundaries and lattice vectors $\bm{L}\in\mathbb{R}^{3\times 3}$. During training, all three components $\left\{\bm{A},\bm{X},\bm{L}\right\}$ are considered simultaneously. We apply the SI framework only to the continuous structural representations $\left\{\bm{X},\bm{L}\right\}$ with loss functions defined in Eqs (2) and (3), and use discrete flow matching (DFM) on the chemical species $\bm{A}$ (see Section III.2.3[43]. The number of atoms $N$ in the structure $x_{0}$ sampled from the base distribution $\rho_{0}$ is determined by the number of atoms in the corresponding structure $x_{1}$ sampled from the target distribution $\rho_{1}$.
@@ III.2.1 Atomic Coordinates
For treating fractional coordinates, we implement a variety of periodic interpolants that connect the base to the target data distributions (see Section IV.1). We specify the base distribution for the fractional coordinates $\bm{x}\in[0,1)$ for all $\bm{x}\in\bm{X}$ via a uniform distribution (except for the score-based diffusion interpolant that, following the approach of Jiao et al. [30], uses a wrapped normal distribution $\rho_{0}(\bm{x})$ which becomes a uniform distribution in the limit of large variance). To satisfy periodic boundary conditions on the paths defined by the interpolants, we extend the SI framework to the surface of a four-dimensional torus in this paper. Reminiscent of RFM [42], the linear interpolant on the torus traverses a path equivalent to the shortest-path geodesic which is always well-defined. [^2] The only exception being when two points are precisely half the box length apart. However, this case is not relevant for the given base distribution. Other interpolants, however, are more complex. In order to uniquely define them, we always define the interpolation with respect to the shortest-path geodesic. That is, for interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with a periodic boundary at 0 and $1$, we first unwrap $\bm{x}_{1}$ to the periodic image $\bm{x}_{1}^{\prime}$ which has the shortest possible distance from $\bm{x}_{0}$. Following this, the interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}^{\prime}$ is computed given a choice of interpolant, and the traversed path is wrapped back into the boundary from 00 to $1$. This approach is illustrated in Appendix B.5.
@@ III.2.2 Lattice Vectors
Lattice vectors $\bm{L}$ are treated with a wide range of (non-periodic) stochastic interpolants (see Section IV.1 again). To construct the base distribution, we follow Miller et al. [32] and construct an informative base distribution $\rho_{0}(\bm{L})$ by combining a uniform distribution over the lattice angles with a log-normal distribution fitted to the empirical distribution of the lattice lengths in each target dataset. This choice brings the base distribution closer to the target distribution. Unlike SBDM, which requires a Gaussian base distribution, the SI framework allows such flexibility. Importantly, the model still has to learn to generate a joint, correlated distribution of lattice vectors, fractional coordinates, and atomic species.
@@ III.2.3 Atomic Species
In doing this, we are able to directly construct the marginal rate matrix $R^{\theta}_{t}(\bm{a}_{t},i)$ for the CTMC that dictates the rate of $\bm{a}_{t}$ at time $t$ jumping to a different state $i$ during generation (see Appendix B.6). It is important to note that the learned probability path is a function of the entire atomic configuration $\left\{\bm{A},\bm{X},\bm{L}\right\}$ which is necessary for the prediction of chemical composition from structure.